# Family Matters: Language Transfer and Merging for Adapting Small LLMs to Faroese

### v1

@@ Abstract
We investigate how to adapt small, efficient LLMs to Faroese, a low-resource North Germanic language. Starting from English models, we continue pre-training on related Scandinavian languages, either individually or combined via merging, before fine-tuning on Faroese. We compare full fine-tuning with parameter-efficient tuning using LoRA, evaluating their impact on both linguistic accuracy and text comprehension. Due to the lack of existing Faroese evaluation data, we construct two new minimal-pair benchmarks from adapted and newly collected datasets and complement them with human evaluations by Faroese linguists. Our results demonstrate that transfer from related languages is crucial, though the optimal source language depends on the task: Icelandic enhances linguistic accuracy, whereas Danish boosts comprehension. Similarly, the choice between full fine-tuning and LoRA is task-dependent: LoRA improves linguistic acceptability and slightly increases human evaluation scores on the base model, while full fine-tuning yields stronger comprehension performance and better preserves model capabilities during downstream fine-tuning.
@@ 1 Introduction
While large language models (LLMs) excel in English and other high-resource languages, low-resource languages lag behind: model quality is tightly linked to data availability Robinson et al. (2023); Li et al. (2024), and even basic comprehension may fail Court and Elsner (2024). Coverage of these languages typically requires the largest available models, if it works at all.
@@
In this paper, we focus on Faroese, a North Germanic language spoken by around 70,000 people, mostly in the Faroe Islands. Training data is scarce, with only 95 million words available in the deduplicated Fineweb-2 dataset Penedo et al. (2024). Faroese is particularly interesting for studying transfer from related languages. In the Nordic context, it has been described as the central Nordic language Torp (1998), reflecting its typological position in relation to the other Scandinavian languages. Derived from Old Norse, shaped by historical development and sociopolitical circumstances, modern Faroese shares characteristics with all the other Scandinavian languages.
@@
We investigate methods for adapting small generative models to Faroese through continued pre-training, with particular emphasis on leveraging transfer from related languages. Prior work has demonstrated that syntactic similarity is a strong predictor of transfer success Chang et al. (2024), and for Faroese, encoder models benefit notably from transfer from Icelandic and other Scandinavian languages Snæbjarnarson et al. (2023). Beyond direct transfer, we explore parameter merging (Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023), which enables the combination of models fine-tuned on different languages. This approach allows us to control the relative influence of each source language, a flexibility that would be more costly to achieve via multilingual training. Our goal is to balance the close linguistic similarity of Icelandic with the larger data resources available for mainland Scandinavian languages. To this end, we merge models trained on various Scandinavian languages and subsequently fine-tune the merged models on Faroese. We also compare two fine-tuning strategies for the continued pre-training: full fine-tuning and parameter-efficient fine-tuning with LoRA (Hu et al., 2022). Prior evidence suggests that full fine-tuning provides higher accuracy, while LoRA better preserves previously learned skills and diversity in text generation (Biderman et al., 2024). Based on this, we hypothesize that full fine-tuning will yield stronger linguistic performance in Faroese, but with a higher risk of losing general reasoning and knowledge abilities.
@@
A major challenge for low-resource languages is the scarcity of evaluation data. To address this, we take two steps. First, we introduce two minimal-pair evaluation suites for small LLMs, combining pre-existing, newly collected, and adapted datasets: FoBLiMP, which targets syntactic and linguistic acceptability, and FoBCoMP, which targets text comprehension. Second, we perform expert human evaluations (carried out by Faroese linguists) on two tasks, one with and one without low-resource downstream task fine-tuning.
Our results indicate that transfer from related languages is crucial. Transfer from Icelandic yields the best performance in the linguistic probes in FoBLiMP, whereas higher-resource mainland Scandinavian languages contributes more to the text comprehension probes in FoBCoMP. Language merging shows promise for combining these advantages, but improvements are not consistent. Contrary to our expectations, LoRA outperforms full fine-tuning on linguistic acceptability in both automatic and human evaluations. In contrast, full fine-tuning achieves better results on comprehension tasks and coupled with downstream fine-tuning.
Our research questions are the following:
@@ 1. RQ1
    What is the effect of transfer from (to various degrees) related languages?
@@ 2. RQ2
    Does merging various related languages have a benefit over choosing the closest neighbor?
@@ 3. RQ3
    What are the differences between fine-tuning all parameters of the base model and performing adaptation with LoRA?
2 Related Work
Language Adaptation
Post-hoc adaptation transfers knowledge from a high-resource or multilingual model to a new language, often using parameter-efficient methods. Early work focused on encoder models (Pfeiffer et al., 2020; Ansell et al., 2022; Ebrahimi and Kann, 2021). Yong et al. (2023) found that full fine-tuning works best for smaller models (e.g., 560M), while adapters are better for larger ones (up to 7.1B), but despite using a generative model, they do not evaluate on generative tasks. Razumovskaia et al. (2024) show that continued pre-training with LoRA (similar to our setup) improves linguistic quality of generations, though not few-shot ability. Instruction tuning on translated data is another option for language adaptation (Muennighoff et al., 2023b), but it is less reliable for small languages, and even correct translations can produce “translationese” (Gellerstam, 1986).
Selecting effective transfer languages is crucial for language adaptation. Chronopoulou et al. (2023) show that sharing adapter parameters across related languages improves machine translation for low-resource languages. Similarly, Faisal and Anastasopoulos (2022) find that linguistically informed adapter designs benefit unseen languages. In machine translation, transfer success often depends more on source corpus size and subword overlap than on broader linguistic similarity (Lin et al., 2019). Pairing a low-resource language with a typologically similar, higher-resource language can outperform using all available languages (Neubig and Hu, 2018). However, they focus on translation from a low-resource language; a setup that does not require generation in the small language.
Faroese Typology and Its Implications
@@ 2 Related Work
Faroese is classified as Insular Scandinavian with Icelandic, while Norwegian, Swedish, and Danish are Mainland Scandinavian. This division reflects differences in lexicon, morphology, and syntax, though Faroese also shares traits with Mainland Scandinavian languages Thráinsson et al. (2012). Lexically, Faroese shares words and cognates with Icelandic and Danish, but no quantitative analysis exists Jacobsen (2021, 2022). Morphologically, it is more similar to Icelandic, with some overlap with Norwegian Torp (1998). Syntactically, Faroese is closer to Mainland Scandinavian languages Ussery and Petersen (2023); Debess (2017); Petersen (2010); Petersen and Heycock (2017); Sandøy (2005). Because of this mixed profile, identifying a single closest neighbour language is difficult. In our experiments, we fine-tune using data from all Scandinavian languages, with particular attention to Icelandic, to capture the different typological influences.
Evaluation for low-resource languages and small models
@@
is challenging because there is little task-specific data. On the other hand, few-shot evaluations often focus on knowledge-intensive tasks or tasks that require structured outputs, making them out of scope for small models. Perplexity is the simplest measure of a model’s fit to a language, and only requires held-out text. In low-resource languages, however, fine-grained tokenization from common tokenizers reduces its usefulness (Oh and Schuler, 2024). An alternative is information parity, which compares the negative log-likelihood of a target-language text to its English version (Tsvetkov and Kipnis, 2024). This metric correlates with downstream performance but requires parallel data. Another option is translated benchmarks, but these miss culture-specific content (Chen et al., 2024), introduce translation errors, and produce “translationese”, making the target language artificially easier for the model.
@@
Minimal pairs consist of two sentences that differ slightly, with one correct and one incorrect. Models are expected to assign higher probability to the correct sentence, capturing implicit preferences (Marvin and Linzen, 2018; He et al., 2025). Classical examples include subject-verb agreement (Linzen et al. (2016): “The key is on the table.” vs. “The key are on the table.”) and negative polarity items (Marvin and Linzen (2018): “No students have ever lived here.” vs. “Most students have ever lived here.”). BLiMP (Warstadt et al., 2020) provides a wide range of syntactic minimal pairs, with multilingual extensions like MultiBLiMP (Jumelet et al., 2025) including Faroese. COMPS (Misra et al., 2023) tests semantic knowledge (e.g., “A robin can fly.” vs. “A penguin/table can fly.”), and its multilingual extension (He et al., 2025) exists, though it does not cover Faroese.
@@
Recent work for Faroese has focused on machine translation Scalvini and Debess (2024); Scalvini et al. (2025a); Simonsen and Einarsson (2024); Debess et al. (2025), combining automatic and human assessments. Automatic metrics such as BLEU or chrF prove shallow and fail to capture Faroese-specific nuances Scalvini et al. (2025b). Embedding-based metrics, such as BERTScore via FoBERT Snæbjarnarson et al. (2023), are emerging but limited and rarely validated against human judgment. Human evaluation remains essential. Beyond MT, small Faroese evaluation datasets exist for sentiment analysis Debess et al. (2024) and question answering Simonsen et al. (2025).
@@ 3 Experimental Setup
In this section, we introduce our experimental setup including model training (Section 3.1), merging (Section 3.2), collection of data for the automatic evaluation (Section 3.3), and the human evaluation process and data (Section 3.4). [^1] We make models and data available in a huggingface collection: huggingface.co/collections/jekunz/faroese-adaptation-68d6aea566b16d4f57180682.
@@ 3.1 Training
We use the two smaller SmolLM2 (Allal et al., 2025) models (135M and 360M parameters) as they are fully open, including their training data, and well-trained for their size on an English corpus. Although their size limits performance on knowledge-intensive tasks, it allows us to continually pre-train and compare them in different setups across substantial corpora. We experiment with two adaptation setups to answer RQ3: full-parameter fine-tuning and LoRA fine-tuning. We train for 5 epochs on the Faroese corpus (following Muennighoff et al. (2023a)’s scaling law for data-constraint training); we do not repeat data for other languages. Training details can be found in Table 6 in Appendix A. We train on the deduplicated Fineweb-2 (Penedo et al., 2024) portions for the Scandinavian languages, containing 27B tokens for Danish, 25B for Swedish, 30B for Norwegian-Bokmål, 1.6B for Icelandic, 495M for Norwegian-Nynorsk, and 95M for Faroese. Due to resource constraints, we limit the corpora for Swedish, Danish and Norwegian (Bokmål) to 4B tokens. We perform sequential continued pre-training for RQ1, first on an individual transfer language, then on the Faroese. We do not merge the data of the source languages with the Faroese data because this may introduce language mixing issues (Li et al., 2025).

### v2

@@ Abstract
We investigate strategies for adapting small, efficient language models to Faroese, a low-resource North Germanic language. Starting from English-pretrained models, we apply continued pre-training on related Scandinavian languages—individually or combined via model merging—before fine-tuning on Faroese. We compare full fine-tuning with parameter-efficient adaptation via LoRA, assessing their effects on general language modeling performance, linguistic accuracy, and text comprehension. To address the lack of existing Faroese evaluation resources, we construct two new minimal-pair probing benchmarks, one for linguistic acceptability and one for text comprehension, and complement them with human evaluations conducted by native Faroese linguists. Our results show that transfer from related languages is essential, but the optimal source language is task-dependent: Icelandic improves linguistic accuracy, while Danish boosts reading comprehension. The choice of adaptation method likewise depends on the target task: LoRA yields stronger linguistic acceptability and marginally higher human evaluation scores, whereas full fine-tuning produces better comprehension performance and more robust downstream fine-tuning. Merging multiple related languages under full fine-tuning (but not LoRA) improves general language modeling, though its benefits in the linguistic acceptability and comprehension probes are less consistent.
@@ 1 Introduction
While large language models (LLMs) excel in English and other high-resource languages, low-resource languages lag behind: model quality is tightly linked to data availability Robinson et al. (2023); Li et al. (2024), and even basic comprehension may fail Court and Elsner (2024). Adequate coverage of these languages typically requires the largest available models, if it works at all.
@@
In this paper, we focus on Faroese, a North Germanic language spoken by approximately 70,000 people, primarily in the Faroe Islands. Training data is scarce, with only 95 million words available in the deduplicated FineWeb-2 dataset Penedo et al. (2024). Faroese is typologically well-suited for studying cross-lingual transfer: descended from Old Norse and shaped by a distinctive historical and sociopolitical trajectory, it has been described as the central Nordic language Torp (1998), sharing structural features with all other Scandinavian languages to varying degrees.
@@
We investigate methods for adapting small generative models to Faroese through continued pre-training, with particular emphasis on transfer from related languages. Prior work has shown that syntactic similarity is a strong predictor of transfer success Chang et al. (2024), and Faroese encoder models benefit notably from Icelandic and other Scandinavian languages Snæbjarnarson et al. (2023). Beyond single-language transfer, we explore parameter merging (Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023) as a means of combining models trained on different languages. This allows fine-grained control over the relative influence of each source language—a flexibility that would be costly to achieve through multilingual training—and enables us to balance the closer morphological and lexical proximity of Icelandic against the syntactic similarity and larger data resources of the mainland Scandinavian languages.
We also compare two adaptation strategies: full fine-tuning and parameter-efficient fine-tuning with LoRA (Hu et al., 2022). Prior work suggests that full fine-tuning yields higher task accuracy, while LoRA better preserves previously acquired skills and generation diversity (Biderman et al., 2024). Our initial hypothesis was therefore that full fine-tuning will produce stronger Faroese linguistic performance, but at a greater risk of degrading general reasoning and knowledge capabilities.
@@
A persistent challenge for low-resource languages is the scarcity of evaluation data. We address this in two ways. First, we introduce two minimal-pair evaluation suites: FoBLiMP, targeting linguistic acceptability, and FoBCoMP, targeting text comprehension, both combining pre-existing, newly collected, and adapted datasets. [^1] Both benchmarks are released on the HuggingFace Hub under identifiers jekunz/FoBLiMP and jekunz/FoBCoMP. Second, we conduct expert human evaluations by native Faroese linguists on two generation tasks, one with and one without downstream fine-tuning.
To summarize, our research questions are:
@@ 1. RQ1
    What is the effect of transfer from languages with varying typological proximity?
@@ 2. RQ2
    Does merging multiple related languages offer advantages over selecting the single closest neighbor?
@@ 3. RQ3
    How does full fine-tuning of all parameters compare to LoRA-based adaptation?
Our results confirm that transfer from related languages is essential. Icelandic transfer yields the strongest performance on linguistic probes, while higher-resource mainland Scandinavian languages contribute more to the comprehension probes. Model merging shows promise as language modeling scores improve for full fine-tuning, though gains are not consistent for the more fine-grained probes. Contrary to our initial hypothesis, LoRA outperforms full fine-tuning on linguistic acceptability in both automatic and human evaluations; full fine-tuning performs better on comprehension tasks and when fine-tuned on text summarization.
2 Background
2.1 Faroese Typology
@@
Faroese belongs to the Insular Scandinavian branch together with Icelandic, while Norwegian, Swedish, and Danish form Mainland Scandinavian. This classification reflects differences in lexicon, morphology, and syntax, although Faroese also shares substantial traits with the Mainland languages Thráinsson et al. (2012). Lexically, Faroese shares cognates with both Icelandic and Danish, though no quantitative study exists Jacobsen (2021, 2022). Morphologically, it is closest to Icelandic, with some overlap with Norwegian Torp (1998). Syntactically, however, Faroese aligns more closely with Mainland Scandinavian languages Ussery and Petersen (2023); Debess (2017); Petersen (2010); Petersen and Heycock (2017); Sandøy (2005). Faroese therefore occupies an intermediate position: morphologically and lexically closest to Icelandic, but syntactically closer to the Mainland languages. This mixed profile complicates the identification of a single closest relative and makes cross-lingual transfer particularly relevant.
To quantify Faroese’s position within the Scandinavian family, we compute lexical and syntactic similarity measures. Faroese is not represented in major typological databases and lacks proper vector-based representations in multilingual toolkits such as lang2vec (Littell et al., 2017; Dryer and Haspelmath, 2013), therefore we rely on alternative resources.
Language Lexical Syntactic
Danish 0.167 0.814
Icelandic 0.321 0.803
Norwegian, Bokmål 0.180 0.812
Norwegian, Nynorsk 0.205 0.790
Swedish 0.181 0.794
Table 1: Lexical and Syntactic Similarity of Faroese to other Scandinavian languages. Lex. Sim.: arithmetic mean of N-gram overlap and normalized Levenshtein similarity (computed on Islex lemma pairs). Synt. Sim.: arithmetic mean of three metrics: POS-tag distribution overlap, POS-tag N-gram overlap, and POS-sequence normalized Levenshtein similarity (computed on Flores-200).
For lexical similarity, we use the pan-Nordic dictionary Islex Úlfarsdóttir (2014). We compute normalized Levenshtein distance and character N-gram overlap between Faroese lemmas and their equivalents in other Scandinavian languages; the overall score is the unweighted mean of these measures; results are given in Table 1. While this metric does not explicitly model cognate relations or orthographic differences, it reproduces the hierarchy suggested in the linguistic literature and confirms Icelandic as lexically closest. For syntactic similarity, we use the parallel splits of Flores-200 Costa-jussà et al. (2022). A total of 2007 parallel sentences were POS-tagged using Stanza Qi et al. (2020). We compute three tag-based metrics: global POS distribution overlap, POS N-gram overlap, and normalized Levenshtein similarity on POS sequences; the final score is their unweighted mean. All languages show high syntactic similarity to Faroese, with Danish being the closest. The results support existing linguistic descriptions: Icelandic is morphologically and lexically closest to Faroese, while Mainland Scandinavian languages show stronger syntactic similarity. This typological profile motivates our experimental design, where we consider all Scandinavian languages, with particular emphasis on Icelandic.
2.2 Language Adaptation
Language adaptation refers to updating a pre-trained model to improve its performance in a target language through continued pre-training on target-language data. Early work focused on encoder models (Pfeiffer et al., 2020; Ansell et al., 2022; Ebrahimi and Kann, 2021). For generative models, Yong et al. (2023) found that full fine-tuning works best for smaller models (e.g., 560M parameters), while adapter-based methods are preferable for larger ones (up to 7.1B); however, despite using a generative model, they do not evaluate on generative tasks. Razumovskaia et al. (2024) show that continued pre-training with LoRA improves the linguistic quality of generations, though not few-shot task performance; a setup closely related to ours. Kunz (2025) test different parameter-efficient methods for language adaptation, finding that LoRA placed in the feed-forward layers was the most promising; a setup which we follow in this work.
While work on parameter-efficient language adaptation is relatively sparse, there is extensive research on parameter-efficient task adaptation and instruction fine-tuning. A recurring theme is the interaction between PEFT and catastrophic forgetting. Whitehouse et al. (2024) show that LoRA outperforms full fine-tuning in low-data and zero-shot cross-lingual settings, where full fine-tuning causes models to lose the ability to generate in non-training languages. Biderman et al. (2024) observe a similar pattern in domain adaptation, and Vu et al. (2022) find that prompt tuning is less prone to defaulting to the source language in zero-shot cross-lingual transfer. Beyond forgetting, Ghosh et al. (2024) report that full fine-tuning leads to greater memorization of the instruction-tuning dataset and thus more hallucination, whereas LoRA encourages generalization from pre-training knowledge. Shuttleworth et al. (2025) further show that LoRA forgets less even when overall performance matches full fine-tuning, indicating that the effect is not simply due to underfitting. However, they also find that with large learning rates, LoRA can sometimes forget more than full fine-tuning. It remains unclear whether these findings extend to language adaptation and how LoRA’s restricted parameter space affects adaptation success.
Selecting effective transfer languages is crucial for language adaptation. Chronopoulou et al. (2023) show that sharing adapter parameters across related languages improves translation for low-resource languages, and Faisal and Anastasopoulos (2022) find that linguistically informed adapter designs benefit unseen languages. In translation, syntactic similarity is predictive of transfer performance (Lee et al., 2022), although transfer success may depend more on source corpus size and subword overlap than on broader linguistic similarity (Lin et al., 2019). Pairing a low-resource language with a typologically similar, higher-resource language can outperform using all available languages (Neubig and Hu, 2018). Notably however, these studies focus on translation from a low-resource language, a setup that does not require generation in the low-resource language itself.
2.3 Merging
Model merging provides an efficient alternative to multi-task training by combining several fine-tuned models into a single checkpoint, often improving generalization. In simple linear merging (Wortsman et al., 2022; Choshen et al., 2022), the parameters of the models are averaged using scalar weights $\alpha_{i}$ that determine how much each model influences the final result. Task Arithmetic (Ilharco et al., 2023) instead operates on task vectors, defined as the difference between a fine-tuned model and a shared base model. These vectors represent the updates introduced by each task. The merged model is formed by adding a weighted combination of these task vectors back to the base model, again with weight coefficients $\alpha_{i}$. TIES (Yadav et al., 2023) further reduces interference between tasks. First, it introduces a density hyperparameter $p\in(0,1]$ that determines how many parameters from each task vector are retained. Second, TIES resolves sign conflicts across models. For each parameter, it determines the dominant update direction across tasks and enforces this consensus direction. The final merged model is obtained by adding the weighted, sparsified, and sign-aligned task vectors back to the base model.
For task fine-tuning, merged models often preserve base model performance and have been shown to generalize better to unseen tasks; merging models fine-tuned on multiple tasks separately matches or exceeds joint multi-task training (Yadav et al., 2024). But as with PEFT methods, the effect of merging languages rather than tasks is less explored. Aakanksha et al. (2024) show that models safety-aligned on one language at a time and then merged outperforms data mixing. For continued pre-training however, Glocker et al. (2025) find that merged models’ performance is lower than the performance of the monolingual base models.
A key challenge in merging LoRA models is parameter interference, where updates from different models conflict in the shared parameter space. Stoica et al. (2024) trace this to LoRAs being less aligned than fully fine-tuned models, and propose a weight alignment procedure as a remedy. Tang et al. (2024) instead propose linearizing the LoRA modules prior to merging, which they show improves post-merge performance. Zhao et al. (2025) address interference by clustering LoRAs by similarity, merging within clusters, and ensembling the results. Zhang and Zhou (2025) take a different approach, constraining LoRA subspaces to be orthogonal during fine-tuning so that the resulting models are more mergeable by construction.
2.4 Low-Resource Evaluation
@@
Evaluation for low-resource languages and small models poses a fundamental challenge: task-specific data is scarce, and few-shot evaluations tend to target knowledge-intensive tasks poorly suited to small models. Perplexity offers the simplest measure of language model fit, requiring only held-out text. However, coarse tokenization inflates scores for low-resource languages in ways that do not reflect true linguistic competence (Oh and Schuler, 2024). Information parity addresses this by comparing the negative log-likelihood of a target-language text to that of its English translation (Tsvetkov and Kipnis, 2024); while this metric correlates with downstream performance, it requires parallel data. Translated benchmarks are another common recourse, but they miss culture-specific content (Chen et al., 2024), introduce translation artifacts, and produce translationese—unnatural target-language text that models may find spuriously easier to process.
@@
Minimal pairs consist of two near-identical sentences, one grammatical and one not, with models expected to assign higher probability to the correct sentence (Marvin and Linzen, 2018; He et al., 2025). Classical examples include subject-verb agreement (Linzen et al. 2016: “The key is | are on the table.”) and negative polarity items (Marvin and Linzen 2018: “No | Most students have ever lived here.”). BLiMP (Warstadt et al., 2020) provides broad syntactic coverage, with its multilingual extension MultiBLiMP (Jumelet et al., 2025) including Faroese. COMPS (Misra et al., 2023) tests semantic knowledge, but its multilingual extension (He et al., 2025) does not cover Faroese.
@@
Recent work on Faroese has focused primarily on machine translation (Scalvini and Debess, 2024; Scalvini et al., 2025a; Simonsen and Einarsson, 2024; Debess et al., 2025), where automatic metrics such as BLEU and chrF fail to capture language-specific nuances (Scalvini et al., 2025b). Embedding-based metrics via FoBERT (Snæbjarnarson et al., 2023) are emerging but rarely validated against human judgment, which remains essential. Beyond translation, small datasets exist for sentiment analysis (Debess et al., 2024) and question answering (Simonsen et al., 2025).
@@ 3.1 Training
We use the two smaller SmolLM2 (Allal et al., 2025) models (135M and 360M parameters) as they are fully open, including their training data, and well-trained for their size on an English corpus. Although their size limits performance on knowledge-intensive tasks, it allows us to continually pre-train and compare them in different setups across substantial corpora. We experiment with two adaptation setups to answer RQ3: full-parameter fine-tuning and LoRA fine-tuning. We train for 5 epochs on the Faroese corpus (following Muennighoff et al. (2023)’s scaling law for data-constraint training); we do not repeat data for other languages. Training details can be found in Table 7 in Appendix A. We train on the deduplicated Fineweb-2 (Penedo et al., 2024) portions for the Scandinavian languages, containing 27B tokens for Danish, 25B for Swedish, 30B for Norwegian-Bokmål, 1.6B for Icelandic, 495M for Norwegian-Nynorsk, and 95M for Faroese. Due to resource constraints, we limit the corpora for Swedish, Danish and Norwegian (Bokmål) to 4B tokens. We perform sequential continued pre-training for RQ1, first on an individual transfer language, then on the Faroese. We do not merge the data of the source languages with the Faroese data because this may introduce language confusion and language mixing in model outputs (Li et al., 2025). [^2] We make models and data available in a HuggingFace collection.